# Age-Gender-Country-Specific Death Rates Modelling and Forecasting: A Linear Mixed-Effects Model

### v1

@@ 1 Introduction
Over the years, statisticians, actuaries, and demographers have been thinking about designing a suitable mortality projection model that can be used in the life insurance business, the pension funds industry, social security, epidemiology, and population studies. In the twentieth century, human mortality rates have declined remarkably around the world, and life expectancy has increased steadily. There is no clear indication that this trend of increasing life expectancy has reached its maximum (Burger et al., 2012). The continuing increase in life expectancy has brought additional stress in support systems for the elderly, such as governments, health care systems, the life insurance business, and the pension funds industry. For the pension funds industry and the insurance business, the pricing and reserving of annuities depend on future ASDRs. The continuous upward trajectory of life expectancy and its profound impact across various sectors, particularly the life insurance industry, highlights the criticality of developing robust models for projecting mortality rates.
Studying actuarial mortality tables, the British actuary Benjamin Gompertz observed that a law of geometrical progression pervades large portions of different mortality tables for humans (Gompertz, 1825). He discovered that the probability of dying among individuals aged 30-80 years increases exponentially with age, and he proposed to represent the hazard rates as
@@
where $\alpha,\beta>0$, see Macdonald et al. (2018).
@@
in which the time-dependent component $k_{t}$ is then forecast as a random walk with drift (Lee and Carter, 1992). This model cannot be fitted as a regular regression because it does not include covariates. The LC model has become the ground for many extensions to forecasting single, and multiple-population mortality rates. Embedding the LC model in a Poisson regression model, Brouhns et al. (2002) proposed an approach for age–gender-specific mortality rates. The method of P-splines was applied to smoothing and forecasting mortality tables by Currie et al. (2004). Including a cohort effect, Renshaw and Haberman (2006) proposed an extension to the LC model. Another well-known technique is the Cairns-Blake-Dowd (Cairns et al., 2006) model (specifically designed for pensioner ages), which introduces a two-factor model for the development of the mortality curve through time. Hyndman and Ullah (2007) used functional data analysis, nonparametric smoothing techniques, and robust statistics to model mortality rates (functional demographic model). A comparative evaluation by D’Amato et al. (2011) established an overall comparison between the LC and the functional demographic model to the Italian mortality data. These methods focus originally on mortality forecasting of a single population.
If scientists have more samples to test in their experiments, it will increase the certainty of their estimates. In the case of multi-population mortality forecasting, exploring additional common effects from other populations, parameter estimation is more reliable. As a result, it improves the forecast accuracy for ASDRs. Modifying the original LC model, Li and Lee (2005) have generated forecasts for multi-population death rates. The augmented common factor (ACF) model assumes that
@@
for populations $i=1,\cdots,M$, age groups $x=0,1,\cdots,\omega$, and time periods $t=0,1,\cdots,T$. Kleinow (2015) has introduced a common age effect model to multiple populations. For constructing a two-population mortality model Li et al. (2015) have introduced a systematic process. Richman and Wüthrich (2021) have proposed a neural network extension of the LC model. It is a fully connected neural network with multiple hidden layers. One of the advantages of the neural network extension is that the forecasts are not derived using time series methods. But, they are obtained implicitly once the neural network has been trained on a training dataset. For a review of the mortality modelling and forecasting methods, see Booth and Tickle (2008).
@@
Linear mixed-effects models play an increasingly important role in applied statistics (see Pinheiro and Bates, 2006, for an overview of the theoretical and computational aspects of LME models). An age-specific death rate is a mortality rate in a particular age group. For each population, ASDRs are collected sequentially in different years. Mortality rates from the same age group are usually more similar than rates from other age groups. That means rates from the same age group are related to one another. Therefore, ASDRs form a longitudinal dataset (i.e. repeated measurements over time (Frees, 2004; Antonio and Beirlant, 2007) where observations within each age group are correlated. LME models offer a flexible and powerful approach for analyzing longitudinal data, such as ASDRs collected over time in different populations. In the context of mortality forecasting, LME models provide several advantages and valuable insights. At their core, LME models combine fixed effects and random effects to capture both population-level trends and individual-level variations within the data. Fixed effects represent the overall relationships between predictors (such as age, gender, cohort, and mortality covariates) and the response variable (log ASDRs). These fixed effects help uncover general patterns and associations between mortality rates and various factors of interest. For example, they can reveal how mortality rates change with age or differ between genders. However, mortality data often exhibit correlations and dependencies within specific groups. For instance, mortality rates within the same age group are typically more similar to each other than to rates from other age groups. LME models address this by incorporating random effects, which account for the correlated nature of observations within each age group. By modelling the individual-level variations within age groups, genders, and countries, LME models provide a more nuanced understanding of the mortality trends. The use of LME models in mortality forecasting allows for improved predictions and insights. By considering both fixed and random effects, these models provide a comprehensive understanding of the underlying factors influencing mortality rates. They can capture not only the overall population-level trends but also the specific variations within different subgroups. Furthermore, LME models account for the longitudinal nature of the data, where observations are collected sequentially over time. This is crucial in mortality forecasting because it allows for the incorporation of temporal dependencies and the exploration of trends and changes in mortality rates over different time periods. Therefore, LME models offer a valuable framework for mortality forecasting by integrating fixed effects and random effects to account for both population-level trends and individual-level variations within age groups, genders, and countries. By capturing correlations within the data and accounting for the longitudinal nature of the observations, LME models provide a more robust and accurate analysis of mortality trends, leading to improved predictions and valuable insights for understanding and forecasting mortality patterns.
@@
Tsutakawa (1988) proposes a mixed model for analyzing geographic variability in mortality rates, specifically focusing on lung cancer in Missouri. In an actuarial context, Antonio et al. (2006) develop a flexible model for loss reserving by employing the theory of linear mixed models. Unlike traditional claims-reserving techniques, the mixed model allows for individual claim predictions with a sound statistical basis, including confidence regions. Antonio and Beirlant (2007) explore the use of generalized linear mixed models (GLMMs) for analyzing repeated measurements and longitudinal data in actuarial statistics. Hughes and Haran (2013) propose a novel parameterization of the spatial generalized linear mixed model to address spatial dependence in non-Gaussian spatial data. Their method effectively tackles spatial confounding and improves computational efficiency by reducing the dimensionality of spatial random effects. Lima et al. (2021) employ a linear mixed effects model to investigate age-specific trends in breast cancer (BC) incidence and mortality rates worldwide from 1990 to 2017. Their objective is to determine whether the increases in BC rates are evident across all age groups and regions and whether these changes can primarily be attributed to decreasing fertility rates. For mortality improvement rates with random effects, Renshaw and Haberman (2021) apply regression approaches within the framework of linear mixed-effects models. Verbeeck et al. (2023) employ linear mixed effects models for the estimation of excess mortality during the COVID-19 pandemic.
@@
In this paper, we concentrate on mixed-effects linear regression. Utilizing classical regression models, such as linear or nonlinear regression, is not appropriate for this longitudinal data, where the subjects are age groups. This is due to the presence of repeated measurements within age groups and potential correlations over time. To adequately account for these dependencies and obtain accurate estimates of variance and p-values, it is necessary to employ a specialized statistical approach like linear mixed effects models. To develop a linear mixed-effects model for mortality rate forecasting, we will utilize two covariates: $k_{ct}$ and $k_{t}$. These covariates play distinct roles in capturing the overall mortality trend movement within our study. Specifically, $k_{ct}$ captures the overall mortality trend across two age groups: ages 00 to $40$ and ages $41$ to $110$, for both females and males within each country at time $t$. In contrast, $k_{t}$ represents the general mortality trend across all age groups for females and males across all countries at time $t$. $k_{ct}$ is calculated by averaging log mortality rates across these two distinct age ranges: ages 00 to $40$ and ages $41$ to $110$. We choose this division to emphasize the specific influence of external factors on mortality rates among different age groups. Utilizing the linear mixed-effects model, we can effectively predict and elucidate the variation in the logarithm of the ASDRs variable, $y_{cgxt}$, based on the covariates $k_{ct}$ and $k_{t}$. The variance in random effect tells us how much variability there is among individuals across ages, and the variance in the residual term describes the amount of variance among observations within each age group. We will see that the sum of the variances of the random effects, $\sigma_{\eta}^{2}$, is higher than the residual variance $\sigma^{2}$. It confirms that observations from the same age group are generally more similar than observations from different age groups. The extent of dependence within the data can be quantified by employing the intra-class correlation coefficient (ICC)
@@
see Hox et al. (2017).
Parameter estimation for the variance components $\sigma_{\eta}^{2}$ and $\sigma^{2}$ is typically done in two steps using a technique called restricted maximum likelihood (REML)(Kenward and Roger, 1997; Pinheiro and Bates, 2006; Jiang and Nguyen, 2021).
@@
A complete picture of the identifiability issues in mortality models is provided by Hunt and Blake (2020). Wang (2013) studies the identifiability of linear mixed-effects models, and gives necessary and sufficient conditions of identifiability. One of the advantages of the LME model is that it is identifiable. We will discuss the identifiabilty of the LME in Section 2.
The paper is arranged as follows. Section 2 defines the linear mixed-effects model of log ASDRs. In Section 3, we apply this LME model to a dataset spanning from 1961 to 2010, comprising twelve populations. We employ a stepwise model selection approach to identify the most suitable model. Subsequently, we extend our analysis to project mortality rates for these twelve populations up to 2019, assessing forecast accuracy by comparing the LME model’s predictions to those of LC models fitted individually to each population. Additionally, we explore life expectancies for these populations, deriving values from the fitted LME model and contrasting them with those derived from LC models. Additionally, we employ both LME and LL models to analyze mortality rates across these populations within age groups spanning from 0 to 100, organized into 5-year age intervals. Furthermore, in Section 3, we delve into an exploration of how we can incorporate GDP per capita in the LME model, recognizing the profound influence of economic factors on mortality patterns. In Section 4, we fit the LME models to fifty-eight single population datasets in the HMD (HMD, 2022) for the observations occurring in the years before 2000, and we forecast mortality rates from the year 2000 to 2019. In Section 5, we employ both the LME and LC model to analyze the present value and solvency capital requirement of a hypothetical portfolio. Section 6 concludes with a critical discussion of the LME model.
@@ 2 The Linear Mixed-Effects ASDRs Model
Projections of future mortality rates and life expectancy are needed to formulate plans for health and social services and economic development. We want to construct models not only for estimating the characteristics of the ASDRs but also for predicting how the mortality rates are most likely to evolve. To do so, we need critical parameters in the model capturing trends and evolution in mortality rates of different age groups over time. The age-specific mortality pattern reflects the underlying death mechanisms and epidemiological characteristics of the age group. Data visualization can often reveal patterns hidden from pure numerical analysis. It is essential to look at some mortality data to identify patterns in the age structure of a population. Figure 1 shows the log ASDRs of twelve populations in $1961-2010$. Each curve is an age-specific death rate over 50 years on a natural logarithm scale. We see a clear evolution of mortality rates going down, which means, on average, people tend to live longer. We also see that the changes over time in mortality rates have been, for most log ASDRs, very regular and relatively slow and stable. We observe that the rates of the same age group in different populations follow a similar pattern. Death rates at some ages change much more rapidly than at other ages. The red curves in the middle of plots in Figure 1 show the most substantial effects, infant mortality rates. Finding the regularity in both age patterns and trends over time, we conclude that the population age structure has a consistent and robust internal pattern.
Figure 1: ASDRs (log) for twelve populations $1961-2010$.
In Figures 2, the thick blue and red curves represent $k_{ct}$, which is calculated by averaging log mortality rates across two distinct age ranges: ages 00 to $40$ and ages $41$ to $110$. We adopt this division to highlight the specific impact of external factors on mortality rates among different age groups. Historical data underscores a striking pattern: mortality rates for individuals in the younger age groups, typically up to $40$ or $50$ years old, exhibit substantial fluctuations during various time periods. This volatility is primarily attributed to external factors and societal shifts, with certain influences, such as the prevalence of AIDS, substance abuse, and increased violence, having a notably more pronounced effect on the younger demographic. These factors have played a significant role in shaping mortality trends (Plat, 2009). Consequently, the introduction of the factor $k_{ct}$ into the LME model is aimed at providing a more comprehensive account of these dynamic influences, particularly their amplified impact on younger age groups.
Figure 2: Plot of $k_{ct}$ for each country in our study ($1961-2010$).
@@
Figure 3: Plot of $k_{t}$: Average of log ASDRs across 12 populations and all age groups ($1961-2010$).
@@ 2.1 The Model Formulation
[…]
Alho (1998) studied neighboring ages for mortality rates and observed correlations of about $0.95$ between them. Lee and Tuljapurkar (1994) say that a perfect correlation can be assumed. Therefore, we can say that $k_{ct}$ and $k_{t}$ are two significant covariates to regress log ASDRs. We propose the development of a linear mixed effects model to capture the relationships between the log ASDRs and the mortality covariates. Within this framework, the mortality covariates are considered as the drivers or exogenous series, while the log ASDRs are treated as dependent variables influenced by these drivers.
@@
$y_{cgx}$ $=\begin{bmatrix}y_{cgx0}\\ y_{cgx1}\\ \vdots\\ y_{cgxT}\end{bmatrix},$ (12)
$X_{cgx}$ $=\begin{bmatrix}1&1&1&k_{0}&k_{c0}&\cdots&k_{0}^{n_{1}}&k_{c0}^{n_{1}}&0-x\\ 1&1&1&k_{1}&k_{c1}&\cdots&k_{1}^{n_{1}}&k_{c1}^{n_{1}}&1-x\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1&1&1&k_{T}&k_{cT}&\cdots&k_{T}^{n_{1}}&k_{cT}^{n_{1}}&T-x\end{bmatrix},$ $\beta$ $=\begin{bmatrix}\beta_{0}\\ \beta_{0x}\\ \beta_{0gx}\\ \beta_{1}\\ \beta_{1gx}\\ \vdots\\ \beta_{{n_{1}}}\\ \beta_{{n_{1}}gx}\\ \gamma_{\rm{cohort}}\end{bmatrix},$ (26)
$Z_{cgx}$ $=\begin{bmatrix}1&k_{0}&k_{0}^{2}&\cdots&k_{0}^{n_{2}}&0-x\\ 1&k_{1}&k_{1}^{2}&\cdots&k_{1}^{n_{2}}&1-x\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 1&k_{T}&k_{T}^{2}&\cdots&k_{T}^{n_{2}}&T-x\end{bmatrix},$ $\eta_{cgx}$ $=\begin{bmatrix}\eta_{0cgx}\\ \eta_{1cgx}\\ \vdots\\ \eta_{n_{2}cgx}\\ \eta_{{\rm{cohort}}_{cgx}}\end{bmatrix},$ (36)
@@
where ${n_{1}}$ and ${n_{2}}$ are non-negative integers; $\Psi$ is a positive-definite, symmetric $(n_{2}+2)\times(n_{2}+2)$ matrix; $\beta$ is the fixed effect vector, and $\eta_{cgx}$ is the random effects vector; $X_{cgx}$ and $Z_{cgx}$ are observed fixed-effects and random-effects regressor matrices; and $\epsilon_{cgx}$ is the within-age-group error vector. We assume that the random effects $\eta_{cgx}$ and the within-age-group errors $\epsilon_{cgx}$ are independent for different age-gender-country groups. Additionally, we assume that they are independent of each other for the same age-gender-country group (see Pinheiro and Bates, 2006, chap. 2).
@@
In the context of linear mixed effects modelling, random effects are systematically defined with an average value of 0. This characteristic signifies that, on average, these random effects do not significantly contribute to the anticipated outcome of the response variable; they are, essentially, clustered around zero. Any structured or non-zero variations observed in the response variable are attributed to the presence of fixed-effects terms within the model. By specifying a mean of 0 for random effects, it implies that, across the entire range of possible random effects within the true population, their combined influence on the response variable averages out to zero. Nevertheless, it is crucial to acknowledge that individual random effects can still take non-zero values. These non-zero values account for the inherent random fluctuations that are not accounted for by the fixed-effects part of the model. This modelling approach enables a distinct separation between systematic influences, effectively captured by the fixed effects, and the unexplained, random variability accounted for by random effects. Therefore, random effects are purposely defined with a mean of 0, and any deviation from this mean within random effects must be included as part of the fixed-effects terms. As a result, if there is a non-zero mean for any term within the random effects, it must be incorporated as a component of the fixed-effects terms. This often leads to a scenario where the columns in the regressor matrix pertaining to random effects form a subset of the columns in the regressor matrix linked to the fixed effects (see Pinheiro and Bates, 2006, chap. 2).
@@
$\eta_{{\rm{cohort}}_{cgx}}$ represents the cohort effect across age, gender, and country. This effect exhibits a modulating structure based on age, gender, and country. The cohort effect represents the impact or influence of the birth cohort to which individuals belong on mortality rates. It suggests that individuals born in different years (cohorts) may experience different mortality rates due to various historical, societal, or environmental factors specific to their birth years. We are inspired by Renshaw and Haberman (2006), who introduced a cohort factor to the Lee-Carter model, along with an age-modulating coefficient. We define the LME model in a way that incorporates this cohort effect, structuring it to account for age, gender, and country influences.

### v2

@@ 1 Introduction
Over the years, statisticians, actuaries, and demographers have been thinking about designing a suitable mortality projection model that can be used in the life insurance business, the pension funds industry, social security, epidemiology, and population studies. In the twentieth century, human mortality rates have declined remarkably around the world, and life expectancy has increased steadily. There is no clear indication that this trend of increasing life expectancy has reached its maximum (burger2012human). The continuing increase in life expectancy has brought additional stress in support systems for the elderly, such as governments, health care systems, the life insurance business, and the pension funds industry. For the pension funds industry and the insurance business, the pricing and reserving of annuities depend on future ASDRs. The continuous upward trajectory of life expectancy and its profound impact across various sectors, particularly the life insurance industry, highlights the criticality of developing robust models for projecting mortality rates.
Studying actuarial mortality tables, the British actuary Benjamin Gompertz observed that a law of geometrical progression pervades large portions of different mortality tables for humans (gompertz1825xxiv). He discovered that the probability of dying among individuals aged 30-80 years increases exponentially with age, and he proposed to represent the hazard rates as
@@
where $\alpha,\beta>0$, see macdonald2018modelling.
@@
in which the time-dependent component $k_{t}$ is then forecast as a random walk with drift (lee1992modeling). This model cannot be fitted as a regular regression because it does not include covariates. The LC model has become the ground for many extensions to forecasting single, and multiple-population mortality rates. Embedding the LC model in a Poisson regression model, brouhns2002poisson proposed an approach for age–gender-specific mortality rates. The method of P-splines was applied to smoothing and forecasting mortality tables by currie2004smoothing. Including a cohort effect, renshaw2006cohort proposed an extension to the LC model. Another well-known technique is the Cairns-Blake-Dowd (cairns2006two) model (specifically designed for pensioner ages), which introduces a two-factor model for the development of the mortality curve through time. hyndman2007robust used functional data analysis, nonparametric smoothing techniques, and robust statistics to model mortality rates (functional demographic model). A comparative evaluation by d2011mortality established an overall comparison between the LC and the functional demographic model to the Italian mortality data. These methods focus originally on mortality forecasting of a single population.
If scientists have more samples to test in their experiments, it will increase the certainty of their estimates. In the case of multi-population mortality forecasting, exploring additional common effects from other populations, parameter estimation is more reliable. As a result, it improves the forecast accuracy for ASDRs. Modifying the original LC model, li2005coherent have generated forecasts for multi-population death rates. The augmented common factor (ACF) model assumes that
@@
for populations $i=1,\cdots,M$, age groups $x=0,1,\cdots,\omega$, and time periods $t=0,1,\cdots,T$. kleinow2015common has introduced a common age effect model to multiple populations. For constructing a two-population mortality model li2015step have introduced a systematic process. richman2021neural have proposed a neural network extension of the LC model. It is a fully connected neural network with multiple hidden layers. One of the advantages of the neural network extension is that the forecasts are not derived using time series methods. But, they are obtained implicitly once the neural network has been trained on a training dataset. For a review of the mortality modelling and forecasting methods, see booth2008mortality.
@@
Linear mixed-effects models play an increasingly important role in applied statistics (see pinheiro2006mixed, for an overview of the theoretical and computational aspects of LME models). An age-specific death rate is a mortality rate in a particular age group. For each population, ASDRs are collected sequentially in different years. Mortality rates from the same age group are usually more similar than rates from other age groups. That means rates from the same age group are related to one another. Therefore, ASDRs form a longitudinal dataset (i.e. repeated measurements over time (frees2004longitudinal; antonio2007actuarial) where observations within each age group are correlated. LME models offer a flexible and powerful approach for analyzing longitudinal data, such as ASDRs collected over time in different populations. In the context of mortality forecasting, LME models provide several advantages and valuable insights. At their core, LME models combine fixed effects and random effects to capture both population-level trends and individual-level variations within the data. Fixed effects represent the overall relationships between predictors (such as age, gender, cohort, and mortality covariates) and the response variable (log ASDRs). These fixed effects help uncover general patterns and associations between mortality rates and various factors of interest. For example, they can reveal how mortality rates change with age or differ between genders. However, mortality data often exhibit correlations and dependencies within specific groups. For instance, mortality rates within the same age group are typically more similar to each other than to rates from other age groups. LME models address this by incorporating random effects, which account for the correlated nature of observations within each age group. By modelling the individual-level variations within age groups, genders, and countries, LME models provide a more nuanced understanding of the mortality trends. The use of LME models in mortality forecasting allows for improved predictions and insights. By considering both fixed and random effects, these models provide a comprehensive understanding of the underlying factors influencing mortality rates. They can capture not only the overall population-level trends but also the specific variations within different subgroups. Furthermore, LME models account for the longitudinal nature of the data, where observations are collected sequentially over time. This is crucial in mortality forecasting because it allows for the incorporation of temporal dependencies and the exploration of trends and changes in mortality rates over different time periods. Therefore, LME models offer a valuable framework for mortality forecasting by integrating fixed effects and random effects to account for both population-level trends and individual-level variations within age groups, genders, and countries. By capturing correlations within the data and accounting for the longitudinal nature of the observations, LME models provide a more robust and accurate analysis of mortality trends, leading to improved predictions and valuable insights for understanding and forecasting mortality patterns.
@@
tsutakawa1988mixed proposes a mixed model for analyzing geographic variability in mortality rates, specifically focusing on lung cancer in Missouri. In an actuarial context, antonio2006lognormal develop a flexible model for loss reserving by employing the theory of linear mixed models. Unlike traditional claims-reserving techniques, the mixed model allows for individual claim predictions with a sound statistical basis, including confidence regions. antonio2007actuarial explore the use of generalized linear mixed models (GLMMs) for analyzing repeated measurements and longitudinal data in actuarial statistics. hughes2013dimension propose a novel parameterization of the spatial generalized linear mixed model to address spatial dependence in non-Gaussian spatial data. Their method effectively tackles spatial confounding and improves computational efficiency by reducing the dimensionality of spatial random effects. lima2021global employ a linear mixed effects model to investigate age-specific trends in breast cancer (BC) incidence and mortality rates worldwide from 1990 to 2017. Their objective is to determine whether the increases in BC rates are evident across all age groups and regions and whether these changes can primarily be attributed to decreasing fertility rates. For mortality improvement rates with random effects, renshaw2021modelling apply regression approaches within the framework of linear mixed-effects models. verbeeck2023linear employ linear mixed effects models for the estimation of excess mortality during the COVID-19 pandemic.
@@
In this paper, we concentrate on mixed-effects linear regression. Utilizing classical regression models, such as linear or nonlinear regression, is not appropriate for this longitudinal data, where the subjects are age groups. This is due to the presence of repeated measurements within age groups and potential correlations over time. To adequately account for these dependencies and obtain accurate estimates of variance and p-values, it is necessary to employ a specialized statistical approach like linear mixed effects models. To develop a linear mixed-effects model for mortality rate forecasting, we will utilize two covariates: $k_{ct}$ and $k_{t}$. These covariates play distinct roles in capturing the overall mortality trend movement within our study. Specifically, $k_{ct}$ captures the overall mortality trend across two age groups: ages 0 to $40$ and ages $41$ to $110$, for both females and males within each country at time $t$. In contrast, $k_{t}$ represents the general mortality trend across all age groups for females and males across all countries at time $t$. $k_{ct}$ is calculated by averaging log mortality rates across these two distinct age ranges: ages 0 to $40$ and ages $41$ to $110$. We choose this division to emphasize the specific influence of external factors on mortality rates among different age groups. Utilizing the linear mixed-effects model, we can effectively predict and elucidate the variation in the logarithm of the ASDRs variable, $y_{cgxt}$, based on the covariates $k_{ct}$ and $k_{t}$. The variance in random effect tells us how much variability there is among individuals across ages, and the variance in the residual term describes the amount of variance among observations within each age group. We will see that the sum of the variances of the random effects, $\sigma_{\eta}^{2}$, is higher than the residual variance $\sigma^{2}$. It confirms that observations from the same age group are generally more similar than observations from different age groups. The extent of dependence within the data can be quantified by employing the intra-class correlation coefficient (ICC)
@@
see hox2017multilevel.
Parameter estimation for the variance components $\sigma_{\eta}^{2}$ and $\sigma^{2}$ is typically done in two steps using a technique called restricted maximum likelihood (REML)(Kenward1997restricted; pinheiro2006mixed; jiang2021linear).
@@
A complete picture of the identifiability issues in mortality models is provided by hunt2020identifiability. wang2013identifiability studies the identifiability of linear mixed-effects models, and gives necessary and sufficient conditions of identifiability. One of the advantages of the LME model is that it is identifiable. We will discuss the identifiabilty of the LME in Section 2.
The paper is arranged as follows. Section 2 defines the linear mixed-effects model of log ASDRs. In Section 3, we apply this LME model to a dataset spanning from 1961 to 2010, comprising twelve populations. We employ a stepwise model selection approach to identify the most suitable model. Subsequently, we extend our analysis to project mortality rates for these twelve populations up to 2019, assessing forecast accuracy by comparing the LME model’s predictions to those of LC models fitted individually to each population. Additionally, we explore life expectancies for these populations, deriving values from the fitted LME model and contrasting them with those derived from LC models. Additionally, we employ both LME and LL models to analyze mortality rates across these populations within age groups spanning from 0 to 100, organized into 5-year age intervals. Furthermore, in Section 3, we delve into an exploration of how we can incorporate GDP per capita in the LME model, recognizing the profound influence of economic factors on mortality patterns. In Section 4, we fit the LME models to fifty-eight single population datasets in the HMD (HMD) for the observations occurring in the years before 2000, and we forecast mortality rates from the year 2000 to 2019. In Section 5, we employ both the LME and LC model to analyze the present value and solvency capital requirement of a hypothetical portfolio. Section 6 concludes with a critical discussion of the LME model.
@@ 2 The Linear Mixed-Effects ASDRs Model
Projections of future mortality rates and life expectancy are needed to formulate plans for health and social services and economic development. We want to construct models not only for estimating the characteristics of the ASDRs but also for predicting how the mortality rates are most likely to evolve. To do so, we need critical parameters in the model capturing trends and evolution in mortality rates of different age groups over time. The age-specific mortality pattern reflects the underlying death mechanisms and epidemiological characteristics of the age group. Data visualization can often reveal patterns hidden from pure numerical analysis. It is essential to look at some mortality data to identify patterns in the age structure of a population. Figure 1 shows the log ASDRs of twelve populations in 196120101961-2010. Each curve is an age-specific death rate over 50 years on a natural logarithm scale. We see a clear evolution of mortality rates going down, which means, on average, people tend to live longer. We also see that the changes over time in mortality rates have been, for most log ASDRs, very regular and relatively slow and stable. We observe that the rates of the same age group in different populations follow a similar pattern. Death rates at some ages change much more rapidly than at other ages. The red curves in the middle of plots in Figure 1 show the most substantial effects, infant mortality rates. Finding the regularity in both age patterns and trends over time, we conclude that the population age structure has a consistent and robust internal pattern.
Figure 1: ASDRs (log) for twelve populations 196120101961-2010.
In Figures 2, the thick blue and red curves represent $k_{ct}$, which is calculated by averaging log mortality rates across two distinct age ranges: ages 0 to $40$ and ages $41$ to $110$. We adopt this division to highlight the specific impact of external factors on mortality rates among different age groups. Historical data underscores a striking pattern: mortality rates for individuals in the younger age groups, typically up to $40$ or $50$ years old, exhibit substantial fluctuations during various time periods. This volatility is primarily attributed to external factors and societal shifts, with certain influences, such as the prevalence of AIDS, substance abuse, and increased violence, having a notably more pronounced effect on the younger demographic. These factors have played a significant role in shaping mortality trends (plat2009stochastic). Consequently, the introduction of the factor $k_{ct}$ into the LME model is aimed at providing a more comprehensive account of these dynamic influences, particularly their amplified impact on younger age groups.
Figure 2: Plot of $k_{ct}$ for each country in our study (196120101961-2010).
@@
Figure 3: Plot of $k_{t}$: Average of log ASDRs across 12 populations and all age groups (196120101961-2010).
@@ 2.1 The Model Formulation
for $c=1,2,\cdots,M$, and $t=0,1,\cdots,T$. $k_{ct}$ and $k_{t}$ will be included as predictors in our model. We will employ random walks with drift to forecast the future values of $k_{ct}$ for ages 0 to $40$ for each country, as well as the future values of $k_{ct}$ for ages $41$ to $\omega$ for each country, and the future values of $k_{t}$. This approach allows us to consider the evolving values and unique temporal patterns exhibited by each country, effectively capturing the specific dynamics and variations in time trends across different countries. We refer to these covariates as the “mortality covariates” due to their significance in capturing the temporal patterns of mortality rates. By incorporating the country mortality covariate, namely $k_{ct}$, into our analysis, we can accurately account for the nuanced temporal dynamics observed within each country, thereby enhancing the accuracy and reliability of our mortality rate pattern analyses. The country mortality covariate play a crucial role in capturing the temporal patterns and enabling a comprehensive understanding of the mortality trends across the countries under study.
Alho1998Stochastic studied neighboring ages for mortality rates and observed correlations of about $0.95$ between them. lee1994stochastic say that a perfect correlation can be assumed. Therefore, we can say that $k_{ct}$ and $k_{t}$ are two significant covariates to regress log ASDRs. We propose the development of a linear mixed effects model to capture the relationships between the log ASDRs and the mortality covariates. Within this framework, the mortality covariates are considered as the drivers or exogenous series, while the log ASDRs are treated as dependent variables influenced by these drivers.
@@
$y_{cgx}$ $=\begin{bmatrix}y_{cgx0}\\ y_{cgx1}\\ \vdots\\ y_{cgxT}\end{bmatrix}$,
$X_{cgx}$ $=\begin{bmatrix}1&1&1&k_{0}&k_{c0}&\cdots&k_{0}^{n_{1}}&k_{c0}^{n_{1}}&0-x\\ 1&1&1&k_{1}&k_{c1}&\cdots&k_{1}^{n_{1}}&k_{c1}^{n_{1}}&1-x\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1&1&1&k_{T}&k_{cT}&\cdots&k_{T}^{n_{1}}&k_{cT}^{n_{1}}&T-x\end{bmatrix}$, $\beta$ $=\begin{bmatrix}\beta_{0}\\ \beta_{0x}\\ \beta_{0gx}\\ \beta_{1}\\ \beta_{1gx}\\ \vdots\\ \beta_{{n_{1}}}\\ \beta_{{n_{1}}gx}\\ \gamma_{\rm{cohort}}\end{bmatrix}$,
@@
where $n_{1}$} and $n_{2}$} are non-negative integers; $\Psi$ is a positive-definite, symmetric $(n_{2}+2)\times(n_{2}+2)$ matrix; $\beta$ is the fixed effect vector, and $\eta_{cgx}$ is the random effects vector; $X_{cgx}$ and $Z_{cgx}$ are observed fixed-effects and random-effects regressor matrices; and $\epsilon_{cgx}$ is the within-age-group error vector. We assume that the random effects $\eta_{cgx}$ and the within-age-group errors $\epsilon_{cgx}$ are independent for different age-gender-country groups. Additionally, we assume that they are independent of each other for the same age-gender-country group (see pinheiro2006mixed, chap. 2).
@@
In the context of linear mixed effects modelling, random effects are systematically defined with an average value of 0. This characteristic signifies that, on average, these random effects do not significantly contribute to the anticipated outcome of the response variable; they are, essentially, clustered around zero. Any structured or non-zero variations observed in the response variable are attributed to the presence of fixed-effects terms within the model. By specifying a mean of 0 for random effects, it implies that, across the entire range of possible random effects within the true population, their combined influence on the response variable averages out to zero. Nevertheless, it is crucial to acknowledge that individual random effects can still take non-zero values. These non-zero values account for the inherent random fluctuations that are not accounted for by the fixed-effects part of the model. This modelling approach enables a distinct separation between systematic influences, effectively captured by the fixed effects, and the unexplained, random variability accounted for by random effects. Therefore, random effects are purposely defined with a mean of 0, and any deviation from this mean within random effects must be included as part of the fixed-effects terms. As a result, if there is a non-zero mean for any term within the random effects, it must be incorporated as a component of the fixed-effects terms. This often leads to a scenario where the columns in the regressor matrix pertaining to random effects form a subset of the columns in the regressor matrix linked to the fixed effects (see pinheiro2006mixed, chap. 2).
@@
$\eta_{{\rm{cohort}}_{cgx}}$ represents the cohort effect across age, gender, and country. This effect exhibits a modulating structure based on age, gender, and country. The cohort effect represents the impact or influence of the birth cohort to which individuals belong on mortality rates. It suggests that individuals born in different years (cohorts) may experience different mortality rates due to various historical, societal, or environmental factors specific to their birth years. We are inspired by renshaw2006cohort, who introduced a cohort factor to the Lee-Carter model, along with an age-modulating coefficient. We define the LME model in a way that incorporates this cohort effect, structuring it to account for age, gender, and country influences.